OpenStudy (anonymous):

1st point: ___6___ 2nd point:___-2___ 3rd point: ___-7___ Using the points above as zeros, construct the polynomial function, f(x), that will be the path of your rollercoaster I need help with this?

OpenStudy (kc_kennylau):

The factor theorem: If \(a\) is a zero of \(f(x)\), then \((x-a)\) is one of its factors.

OpenStudy (kc_kennylau):

Maybe you should find out the factors of the polynomial function.

OpenStudy (anonymous):

How do I turn these numbers into a polynomial function?

OpenStudy (kc_kennylau):

You multiply all its factors together, and expand.

OpenStudy (anonymous):

so multiply 6 -2 and -7??

OpenStudy (kc_kennylau):

Nope, 6, -2 and -7 are not factors of f(x)

OpenStudy (kc_kennylau):

Please kindly read what I have written above

OpenStudy (anonymous):

oh lol sorry! can you tell me how i find the factors?

OpenStudy (kc_kennylau):

The factor theorem: If \(a\) is a zero of \(f(x)\), then \((x-a)\) is one of its factors.

OpenStudy (kc_kennylau):

For example, if 3 were one of its zeros, then (x-3) would be one of its factors.

OpenStudy (anonymous):

I'm so confused lol I'm sorry

OpenStudy (anonymous):

Can you just tell me what to do from the start and maybe I'll understand in the process

OpenStudy (kc_kennylau):

So you know that 6 is one of its zeros... what's its corresponding factor?

OpenStudy (kc_kennylau):

I just said that if 3 were one of its zeros, then (x-3) would be one of its factors

OpenStudy (anonymous):

so it would be (x-6)?

OpenStudy (kc_kennylau):

Exactly

OpenStudy (kc_kennylau):

So what's the corresponding factors for -2 and -7?

OpenStudy (anonymous):

x--2 and x--7?

OpenStudy (anonymous):

which turn into positives right?

OpenStudy (kc_kennylau):

Yep, so you're getting it :)

OpenStudy (anonymous):

so what do i do after i turn them into factors?>

OpenStudy (kc_kennylau):

What are the factors again?

OpenStudy (anonymous):

thank you so much for helping right now lol i'm failing my math class right now and im in desperate need of help lol

OpenStudy (anonymous):

(x-6) (x+3) (x+7)

OpenStudy (kc_kennylau):

One tiny typo out there, spot it :p

OpenStudy (anonymous):

(x+2) lol sorry

OpenStudy (kc_kennylau):

Now expand it :)

OpenStudy (anonymous):

what do you mean?

OpenStudy (kc_kennylau):

Expand (x-6)(x+2)(x+7)

OpenStudy (anonymous):

how do i expand it

OpenStudy (kc_kennylau):

Well...

OpenStudy (kc_kennylau):

FOIL it?

OpenStudy (anonymous):

do i multiply x by x?

OpenStudy (kc_kennylau):

Well... talking about expansion...

OpenStudy (kc_kennylau):

\[(x-6)(x+2)(x+7)\]\[=[x(x+2)-6(x+2)](x+7)\]\[=(x^2+2x-6x-12)(x+7)\]\[=(x^2-4x-12)(x+7)\]\[=x(x^2-4x-12)+7(x^2-4x-12)\]\[=x^3-4x^2-12x+7x^2-28x-84\]\[=x^3+3x^2-40x-84\]

OpenStudy (kc_kennylau):

Any step you don't understand?

OpenStudy (anonymous):

I think that I get it

OpenStudy (kc_kennylau):

Glad to hear that

OpenStudy (anonymous):

so is that all i do for this question?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

thanks so much! do you think you can help me with the next question?

OpenStudy (anonymous):

let me write all of this down first though so i know what i'm doing lol.

OpenStudy (anonymous):

Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph

OpenStudy (kc_kennylau):

Where is the graph?

OpenStudy (anonymous):

oh sorry hold on.

OpenStudy (kc_kennylau):

That ain't a graph?

OpenStudy (anonymous):

that's whats with the assignment lol

OpenStudy (anonymous):

then can you help[ with this one ts my last one? jSolve for the y–intercept for your function, f(x), and then construct a rough graph of your rollercoaster. If your y–intercept is off the graph, give the coordinates of the y–intercept.

OpenStudy (anonymous):

@kc_kennylau